# Knowledge-guided machine learning for county-level corn yield prediction under drought

### v2

@@ 2.2 APSIM field-level dataset
In this section, we explain how APSIM was used to generate a field-level simulated dataset for model pretraining. We used data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides U.S. agricultural and environmental data. We extracted all available station data from the IEM, which included data from thousands of stations across 12 states from 1980 to 2023 (Table 2).
Table 2: The number of stations selected for each state.
@@
Station number 120 89 113 150 140 140 130 128 79 100 114 147
@@ 2.2.1 Input and output of APSIM simulation
The APSIM model simulates corn yield based on four weather data inputs: maximum and minimum temperature (Tmax and Tmin), precipitation (PPT), and radiation (Radn). The simulation also requires the station’s latitude, longitude, and year. Some management parameters are also necessary, including start and end of sowing window, plant population, fertilizer amount, and initial soil water, etc (Sec. 3.3.1). Tmax and Tmin influence the rate of plant development, and affect processes like photosynthesis and respiration. PPT are essential for modeling soil moisture levels, which directly affect water availability for crops. Radn is a critical factor in photosynthesis, as it provides the energy needed for plant growth. These weather data of IEM stations are available from the IEM website (Herzmann et al., 2004). The APSIM model takes weather data as input to simulate root zone soil moisture (SM_rootzone), surface soil moisture (SM_surface), and corn yield. SM_rootzone is crucial for corn growth as it directly affects water availability for uptake, influencing plant development and yield. SM_surface plays a key role in seed germination and early growth stages.
@@ 2.2.2 Variables summary in field-level dataset for pretraining
We calculated a five-year historical average yield based on the simulated corn yield to establish a local baseline for model accuracy improvement. We also included the prediction year and location (longitude and latitude) of IEM stations in the field-level dataset. Because soil properties and topography vary across space and time, year and location can help the ML model capture this variation. Then we combined the weather data, simulated soil moisture and corn yield, along with other variables to construct the APSIM field-level dataset (Table 3).
Table 3: Summary of variables comprising the APSIM field-level dataset, including inputs and simulated outputs
Category Variables
Weather data Radn (MJ/m2) (input)
Tmax (${}^{\circ}\mathrm{C}$) (input)
Tmin (${}^{\circ}\mathrm{C}$) (input)
PPT (mm) (input)
Soil moisture SM_surface (simulated)
SM_rootzone (simulated)
Corn yield Corn yield (t/ha) (simulated)
Others Prediction year (input)
Location (Latitude and longitude) (input)
Historical average yield (t/ha) (simulated)
@@ 2.3 GEE county-level dataset
In this section, we introduce how to construct a county-level dataset for finetuning. Details of the studied counties are provided in Table 4. This dataset includes all the variable types in the APSIM field-level dataset (Table 3), including weather, soil moisture, and other features, but from different sources. To enhance ML model performance with RS data, we also included four Vegetation Indices (VIs) from satellite imagery: Green Chlorophyll Index (GCI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI).
Table 4: The average number of counties selected for each state in GEE county-level dataset.
State IL IN IA KS MI MN MO NE ND OH SD WI
County number 91 74 87 74 57 64 74 69 48 72 39 62
@@ 2.3.1 Vegetation indices
The Terra and Aqua combined Moderate Resolution Imaging Spectroradiometer (MODIS) Land Cover Climate Modeling Grid Version 6 data product (MCD12Q1 v6) (Schaaf and Wang, 2015) is a vital source of satellite-derived data with 500 m resolution, providing consistent, high-quality observations of the Earth’s surface. MODIS MCD12Q1 v6 captures data in multiple spectral bands and offers various products like VIs. These indices are essential for monitoring vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales. They help improve yield prediction by indicating vegetation growth and crop health.
The Green Chlorophyll Index (GCI) (Gitelson et al., 2005) is used to estimate the chlorophyll content in plants, which directly relates to plant health and productivity. It is calculated using the near-infrared (NIR) and green spectral bands. GCI is valuable for assessing vegetation health, especially for monitoring crop growth and detecting areas with nutrient deficiencies. High chlorophyll content typically indicates a healthy, productive plant, making GCI an essential tool in agricultural management. GCI is defined as:
$GCI=\frac{NIR}{Green}-1$ (1)
where $NIR$ represents the surface reflectance in the Near Infrared band, and $Green$ represents the surface reflectance in the Green band.
The Enhanced Vegetation Index (EVI) (Huete et al., 2002) is designed to optimize the vegetation signal by enhancing sensitivity in high biomass areas and reducing atmospheric and canopy background noise. It uses the red, blue, and NIR spectral bands. EVI is particularly useful in regions with dense vegetation, where NDVI might saturate. It provides a more accurate representation of vegetation conditions, especially in areas with high leaf area index, making it ideal for monitoring forests and agricultural lands. EVI is defined as:
$EVI=2.5\times\frac{NIR-Red}{NIR+6\times Red-7.5\times Blue+1}$ (2)
where $Red$ and $Blue$ represent the surface reflectances of the Red band and Blue band, respectively.
The Normalized Difference Water Index (NDWI) (Gao, 1996) measures the moisture content in vegetation and soil by using the NIR and shortwave infrared (SWIR) bands. NDWI is crucial for monitoring water stress in plants, assessing drought conditions, and detecting changes in water bodies. It aids in managing irrigation practices and understanding the impact of water availability on crop yields. NDWI is defined as:
$NDWI=\frac{NIR-SWIR}{NIR+SWIR}$ (3)
where $SWIR$ represents the shortwave infrared band.
The Normalized Difference Vegetation Index (NDVI) (KRIEGLER, 1969) is one of the most widely used indices for measuring vegetation greenness and biomass. It is calculated using the NIR and red spectral bands. NDVI is a standard tool for monitoring plant growth, assessing vegetation cover, and detecting changes in biomass over time. It is particularly useful for tracking seasonal variations in vegetation and estimating crop yields. NDVI is defined as:
$NDVI=\frac{NIR-RED}{NIR+RED}$ (4)
@@ 2.3.2 Weather data
The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product. The reliable radiation data (Radn) is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours. Downward Shortwave Radiation is incident solar radiation over land surfaces in the shortwave spectrum (300-4,000 nanometers) and is an important variable in land-surface models that address a variety of scientific and applied issues.
@@ 2.3.3 Soil moisture
The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Reichle et al., 2022) plays a critical role in our research on the relationship between drought and corn yield prediction. By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP level 4 data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
@@ 2.3.4 Variables summary in county-level dataset for finetuning
Table 5: Summary of variables in GEE county-level dataset
Category Variables Spatial resolution Source
Vegetation index Green Chlorophyll Index (GCI) 500m MODIS
Enhanced Vegetation Index (EVI)
Normalized Difference Water Index (NDWI)
Normalized Difference Vegetation Index (NDVI)
Weather data Radn (W/$m^{2}$)
Tmax (${}^{\circ}\mathrm{C}$) 4km PRISM
Tmin (${}^{\circ}\mathrm{C}$)
PPT (mm)
Soil moisture SM_surface 9km SMAP
SM_rootzone
Corn yield USDA NASS corn yield (t/ha) N/A USDA NASS
Others Prediction year
Location (Latitude and longitude)
Historical average yield (t/ha)
@@ 3 Methodology
We begin with a formal problem formulation of KGML-SM (Sec. 3.1), followed by an introduction to the model components (Sec. 3.2); next, we explain how the APSIM field-level dataset is generated for pretraining (Sec. 3.3) and the GEE county-level dataset for finetuning and testing (Sec. 3.4); finally, we introduce how the KGML-SM model is trained and used for prediction (Sec. 3.5).
@@ 3.2.1 Weather-to-Soil encoder
The W2S encoder is a module designed to model the influence of weather conditions on soil moisture. By capturing the relationship between weather and soil moisture, the W2S encoder improves the model’s understanding of soil dynamics and the impact of weather variability on soil conditions and corn yield.
@@ 3.3.1 APSIM management parameters
Simulation methodology depends on the study’s scale. For field-level areas, management information is obtained directly from farmers as input. Other unknown parameters are adjusted to optimize the model (Shahhosseini et al., 2021). For county-level areas, identifying representative fields is common practice. Their management parameters are averaged, and then applied to represent the entire county (Puntel et al., 2016). Calibrating each county individually became impractical for very large study areas like our U.S. Corn Belt simulation. In such cases, the typical approach defines a parameter range based on empirical and statistical data. These parameters are then randomly combined within specified ranges, simulating yield across all counties (Lobell et al., 2015). Given the vast area of cornfields in our study, it was not feasible to calibrate these parameters for every individual station. Therefore, we established a general parameter range encompassing all possible values across different fields. This approach ensured model applicability across a wide range of conditions, maintaining reasonable prediction accuracy.
@@
Some reasonable adjustment ranges for key management parameters were collected from prior research and USDA statistics (Table 6). Sowing density determines the number of plants per unit area, directly affecting competition for resources such as light, water, and nutrients. Based on statistics from the USDA NASS (USDA, 2020), the range for sowing density was set at 6-9 plants/m2, values of 6,7,8, and 9 were used in the simulation. Sowing dates are crucial as they determine the crop’s growth cycle and its interaction with seasonal weather patterns. Typically, growers maximize corn yield by planting in late April or early May (Licht, 2021; Coulter, 2024). In the simulation, sowing started between April 20–25 and ended between May 15–20. Fertilizer application is essential for providing the necessary nutrients to support plant growth. The most commonly used nitrogen fertilizers for corn production in North America are anhydrous ammonia, urea, and urea-ammonium nitrate solutions (Herzmann et al., 2004). The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N), values of 200, 250, and 300 were used in the simulation. Initial soil water content is important for establishing the starting conditions for the model’s simulation of soil moisture dynamics throughout the growing season. The initial soil water content was set between 40% and 60%, values of 40%, 50%, and 60% were used in the simulation.
@@ 3.3.2 Optimizing APSIM field-level dataset based on soil moisture
APSIM is not accurate for large-scale corn yield simulation without precise management adjustments, thus requiring optimization of the APSIM field-level pretrain dataset. Soil moisture was proposed as a benchmark for filtering high-quality data. First, soil moisture dynamics are relatively simple, strongly correlated with precipitation. In contrast, corn yield is influenced by multiple factors and complex growth processes, hindering accurate modeling in process-based model. Second, simulation processes inherently exhibit cumulative error. Due to the relative simplicity of soil moisture modeling, this error is comparatively smaller. Therefore, the APSIM field-level dataset was optimized based on the quality of soil moisture data. Specifically, an LR model was trained on the GEE county-level dataset. It utilized four weather inputs (Tmax, Tmin, PPT, Radn) to predict soil moisture output. The trained model subsequently predicted soil moisture for the APSIM field-level dataset using its weather data. This prediction was compared against the simulated soil moisture. Simulated soil moisture with a mean squared error (MSE) exceeding 0.5 were discarded ensuring dataset quality.
@@ 3.4.1 Feature Extraction Within Cornfield
To focus on the corn portion of the acquired feature data from GEE, the cornfield areas had to be identified. The Cropland Data Layer (CDL) cropland mask (USDA-NASS, 2017) was utilized extracting feature data corresponding to cropland. The Cropland Data Layer is an annual raster-based dataset. It provides crop-specific land cover information produced by the USDA. This dataset has a 30-meter resolution. The CDL was employed to generate a cornfield mask in GEE, supporting our corn yield prediction study in twelve U.S. Corn Belt states.
@@ 3.4.2 Data preprocessing
County-level features were aggregated into a single value every 16 days, representing agricultural conditions for that period. This process resulted in a data vector for each feature annually for each county. This aggregation method followed the experimental setup from previous research (Ma et al., 2021b; Wang et al., 2025).

### v3

@@ 2.2 APSIM field-level dataset
In this section, we explain how APSIM was used to generate a field-level simulated dataset for model pretraining. We used daily weather data from the Iowa Environmental Mesonet (IEM) (Herzmann et al., 2004), a platform developed by Iowa State University that provides agricultural and environmental observations across the United States. Specifically, we extracted all available station-based records, which include variables such as temperature, PPT, wind, and solar radiation, together with the latitude and longitude of each station. This APSIM field-level dataset covers several thousand stations across 12 states from 1980 to 2023 (Table 2).
Table 2: The number of stations used for APSIM field-level dataset and the average number of counties used for GEE county-level dataset in each state.
@@
Station number
(APSIM field-level dataset)
120 89 113 150 140 140 130 128 79 100 114 147
County number
(GEE county-level dataset)
91 74 87 74 57 64 74 69 48 72 39 62
@@ 2.2.1 Input and output of APSIM simulation
The APSIM model simulates corn yield based on four weather data inputs: maximum and minimum temperature (Tmax and Tmin), PPT, and radiation (Radn). Tmax and Tmin influence the rate of plant development, and affect processes like photosynthesis and respiration. PPT are essential for modeling soil moisture levels, which directly affect water availability for crops. Radn is a critical factor in photosynthesis, as it provides the energy needed for plant growth. These weather data of IEM stations are available from the IEM website (Herzmann et al., 2004). The simulation also requires the station’s latitude, longitude, and year. Some management parameters are also necessary, such as the start and end of the sowing window, plant population, fertilizer amount, and initial soil water, which were described in detail in Sec. 3.3.1. The APSIM model takes weather data as input to simulate root zone soil moisture (SM_rootzone), surface soil moisture (SM_surface), and corn yield. SM_rootzone is crucial for corn growth as it directly affects water availability for uptake, influencing plant development and yield. SM_surface plays a key role in seed germination and early growth stages.
@@ 2.2.2 Variables summary in field-level dataset for pretraining
We adopted a five-year averaging strategy consistent with prior studies (Ma et al., 2021a; Wang et al., 2025), which enhances robustness by mitigating short-term variability while retaining long-term soil–climate signals. The resulting five-year historical average of simulated corn yield (Historical avg. yield) was used as a local baseline to improve model accuracy. We also included the prediction year and location (Lat&Long) of IEM stations in the field-level dataset. Because soil properties and topography vary across space and time, year and location can help the ML model capture this variation. We then combined the weather data, simulated soil moisture and corn yield, along with other variables to construct the APSIM field-level dataset (Table 3). VIs were not included in the pretraining dataset because the IEM stations, although co-located with RS imagery, are not necessarily situated in cornfields, making such signals unreliable for yield modeling. To avoid introducing noise, VIs were set to zero during pretraining, while in the finetuning stage, actual RS-based VIs were incorporated at the county level where crop type and coverage are reliably defined.
Table 3: Summary of variables in the APSIM field-level dataset and the GEE county-level dataset
Category Variables APSIM field-level dataset GEE county-level dataset
Unit Type Source Unit
Spatial
resolution
Source
Weather data Radn MJ/m2 Input IEM W/m2 4 km PRISM
Tmax C Input IEM C 4 km PRISM
Tmin C Input IEM C 4 km PRISM
PPT $m$ Input IEM $m$ 4 km PRISM
Soil moisture SM_surface Simulated APSIM 9 km SMAP
SM_rootzone Simulated APSIM 9 km SMAP
Corn yield Yield t/ha Simulated APSIM t/ha USDA NASS
VIs GCVI 500 m MODIS
EVI 500 m MODIS
NDWI 500 m MODIS
NDVI 500 m MODIS
Others Prediction year Input IEM NASS
Location (Lat&Long) Lat&Long Input IEM Lat&Long USDA NASS
Historical avg. yield t/ha Simulated APSIM t/ha USDA NASS
@@ 2.3 GEE county-level dataset
In this section, the construction of a county-level dataset for finetuning is described. Details of the studied counties are provided in Table 2. This dataset contained all variable types from the APSIM field-level dataset (Table 3), including weather, soil moisture, and other features, but here they were obtained from RS and reanalysis products available in GEE. To enhance our model performance with RS data, four VIs from satellite imagery were also included: Green Chlorophyll Index (GCVI), Enhanced Vegetation Index (EVI), Normalized Difference Water Index (NDWI), and Normalized Difference Vegetation Index (NDVI). Detailed descriptions of these data sources and processing steps are provided in Sec. 2.3.1. Here, we only introduce the variables and data sources of the dataset; the processing methodology is described in Sec. 3.4.
@@ 2.3.1 Vegetation indices
The Terra and Aqua combined Moderate Resolution Imaging Spectroradiometer (MODIS) Land Cover Climate Modeling Grid Version 6 product (MCD12Q1 v6) (Schaaf and Wang, 2015) provides satellite-derived data at 500 m resolution and offers consistent, high-quality observations of the Earth’s surface. MODIS captures information across multiple spectral bands and enables the derivation of various VIs. These indices have been widely used to monitor vegetation health, biomass, water content, and chlorophyll levels across different regions and time scales, thereby improving yield prediction by reflecting crop growth and condition. The VIs used in this study are summarized in Table 4 along with their equations and main applications.
Table 4: Summary of VIs used in this study, including definitions and applications. Abbreviations: NIRNIR (near-infrared reflectance); RedRed (red reflectance); BlueBlue (blue reflectance); GreenGreen (green reflectance); SWIRSWIR (shortwave infrared reflectance). Band specifications for MODIS are as follows: Red – Band 1 (620–670 nm); NIR – Band 2 (841–876 nm); Blue – Band 3 (459–479 nm); Green – Band 4 (545–565 nm); SWIR – Band 6 (1628–1652 nm).
Vegetation Index Citation Equation Application
GCVI (Gitelson et al., 2005) $GCVI=\frac{NIR}{Green}-1$ Chlorophyll content; Crop health; Nutrient status
EVI (Huete et al., 2002) $EVI=\frac{2.5\times(NIR-Red)}{NIR+6\times Red-7.5\times Blue+1}$ Dense canopy; High biomass; Noise reduction
NDWI (Gao, 1996) $NDWI=\frac{NIR-SWIR}{NIR+SWIR}$ Soil/vegetation moisture; Drought monitoring; Irrigation management
NDVI (KRIEGLER, 1969) $NDVI=\frac{NIR-Red}{NIR+Red}$ Vegetation greenness; Biomass; Yield prediction
@@ 2.3.2 Weather data
The MCD18A1 Version 6.1 (Wang, 2021) is a MODIS Terra and Aqua combined Downward Shortwave Radiation gridded Level 3 product. The reliable radiation data is produced daily at 500 m resolution, with estimates of Downward Shortwave Radiation provided every 3 hours. Downward Shortwave Radiation is incident solar radiation over land surfaces in the shortwave spectrum (300-4,000 nanometers) and is an important variable in land-surface models that address a variety of scientific and applied issues.
@@ 2.3.3 Soil moisture
The SPL4SMGP.007 SMAP L4 Global 3-hourly 9-km Surface and Root Zone Soil Moisture dataset (Reichle et al., 2022; Entekhabi et al., 2010) plays a critical role in our research on the relationship between drought and corn yield prediction. By providing detailed measurements of soil moisture at both the surface (0–5 cm) and root zone levels (0–100 cm), SMAP level 4 data allows us to assess the availability of water in the soil, a key factor influencing crop growth and resilience during drought conditions.
@@ 3 Methodology
We begin with a formal problem formulation of KGML-SM (Sec. 3.1), followed by an introduction to the model components (Sec. 3.2); next, we explain how the APSIM field-level dataset is generated for pretraining (Sec. 3.3) and the GEE county-level dataset for finetuning and testing (Sec. 3.4); then, we introduce how the KGML-SM model is trained and used for prediction (Sec. 3.5); after that, we describe the experimental setup (Sec. 3.6) and the statistical analysis of drought, soil moisture, and corn yield (Sec. 3.7).
@@ 3.2.1 Weather-to-Soil encoder
The W2S encoder is a module designed to model the influence of weather conditions on soil moisture. By capturing the statistical relationship between weather variables and soil moisture, the W2S encoder improves the representation of soil moisture dynamics at the county scale, which in turn supports more accurate yield prediction under varying weather conditions.
@@ 3.3.1 APSIM management parameters
Simulation methodology depended on the study’s scale. For field-level areas, management information was obtained directly from farmers as input. Other parameters—such as sowing density, fertilizer amount, initial soil water content, and selected cultivar coefficients—were calibrated to optimize model performance (Zhen et al., 2022). For county-level areas, identifying representative fields was common practice. Their management parameters were averaged, and then applied to represent the entire county (Puntel et al., 2016). Calibrating each county individually became impractical for very large study areas like our U.S. Corn Belt simulation. In such cases, the typical approach defined a parameter range based on empirical and statistical data. These parameters were then randomly combined within specified ranges, simulating yield across all counties (Lobell et al., 2015). Given the vast area of cornfields in our study, it was not feasible to calibrate these parameters for every individual station. Rather than calibrating parameters to specific local conditions, the goal of generating simulated data was to expand the range of training samples, thereby improving the model’s capacity to generalize across diverse environments. Although this dataset may appear relatively coarse due to the lack of calibration, such diversity is in fact desirable for pretraining, as it exposes the model to a wider spectrum of conditions and enhances its ability to generalize. Importantly, our focus is not on accurately predicting yield at these station sites, but rather on ensuring robust county-level performance after finetuning with observed data. At this stage, the simulated dataset primarily serves to expose the model to general patterns of corn growth, while subsequent county-level calibration ensures predictive accuracy at the target scale. By defining parameters within broad but reasonable ranges, we were able to generate a diverse set of simulated samples, which serves as the foundation for the parameter specifications described in the following section.
@@
Therefore, we established a general parameter range encompassing all possible values across different fields (Table 5). This approach ensured model applicability across a wide range of conditions, maintaining reasonable prediction accuracy. Some reasonable adjustment ranges for key management parameters were collected from prior research and USDA statistics. Sowing density determines the number of plants per unit area, directly affecting competition for resources such as light, water, and nutrients. Based on statistics from the USDA NASS (USDA, 2020), the range for sowing density was set at 6-9 plants/m2, values of 6, 7, 8, and 9 were used in the simulation. Sowing dates are crucial as they determine the crop’s growth cycle and its interaction with seasonal weather patterns. Typically, growers maximize corn yield by planting in late April or early May (Licht, 2021; Coulter, 2024). In the simulation, sowing started between April 20–25 and ended between May 15–20. Fertilizer application is essential for providing the necessary nutrients to support plant growth. The most commonly used nitrogen fertilizers for corn production in North America are anhydrous ammonia, urea, and urea-ammonium nitrate solutions (Herzmann et al., 2004). The fertilizer amount was set at 200-300 kg/ha of urea nitrogen (N). Values of 200, 250, and 300 were used in the simulation. Initial soil water content is important for establishing the starting conditions for the model’s simulation of soil moisture dynamics throughout the growing season. The initial soil water content was set between 40% and 60%, values of 40%, 50%, and 60% were used in the simulation.
@@ 3.3.2 Optimizing APSIM field-level dataset based on soil moisture
APSIM was not accurate for large-scale corn yield simulation without precise management adjustments, making it necessary to filter the APSIM field-level dataset before using it for pretraining. Although the goal was to construct a diverse dataset, retaining highly unrealistic samples would only introduce noise and bias the model away from biologically plausible relationships. Filtering was therefore essential to ensure that the dataset remained both diverse and reliable.
Soil moisture was selected as the benchmark variable because it provides a balance between data availability and process relevance. Although it is influenced by multiple factors such as PPT, soil texture, and topography, detailed station-level information on these factors was unavailable across the study region. In contrast, weather data were consistently available and strongly correlated with soil moisture, making it a practical proxy for dataset screening. Moreover, cumulative simulation errors are typically smaller for soil moisture than for yield, which is affected by many interacting processes. Based on these considerations, we used soil moisture quality as the criterion to optimize the APSIM field-level dataset.
Specifically, a Linear Regression (LR) model was trained on the GEE county-level dataset to predict soil moisture from four weather inputs (Tmax, Tmin, PPT, Radn). The trained model then predicted soil moisture predictions for the APSIM field-level dataset, which were compared against the APSIM-simulated values. To ensure data quality while retaining sufficient variability, samples with mean squared error (MSE) above 0.5 were discarded. This threshold was chosen as a balance: higher-error samples (e.g., $\geq$1) were unreliable compared to typical soil moisture RMSE values, yet a more stringent cutoff could have removed too much variability and reduced the generalization capacity of the pretraining dataset.
@@ 3.4.1 Feature extraction within corn field
To focus on the corn portion of the acquired feature data from GEE, the cornfield areas had to be identified. The Cropland Data Layer (CDL) cropland mask (USDA-NASS, 2017) was used to extract the corn class for each year from 2015 to 2023. The CDL is an annual raster-based dataset with a 30-meter resolution, providing crop-specific land cover information produced by the USDA. These annual cornfield masks were employed in GEE to support our corn yield prediction study in twelve U.S. Corn Belt states.
@@ 3.4.2 Data preprocessing
For each county, pixel-level feature values from MODIS and SMAP products were first averaged spatially to obtain county-level values. These county-level values were then aggregated into a single representative mean value every 16 days during the growing season (April–October), following the experimental setup of previous research (Ma et al., 2021b; Wang et al., 2025). This temporal aggregation balanced data availability with noise reduction and provided a consistent seasonal trajectory of vegetation and environmental conditions for each county. We noted that MODIS standard products included atmospheric correction and provided quality assurance layers that flag cloud-affected or low-quality pixels. In this study, we did not apply additional cloud filtering beyond the standard MODIS product processing; instead, the 16-day temporal aggregation helps to smooth residual noise caused by occasional cloud contamination. This procedure resulted in an annual time-series vector of features for each county, which served as input for model training and evaluation.